OpenStudy (anonymous):

Create an addition word problem and a subtraction word problem using fractions. Show how to find the solution to each problem. Once you find the solutions to your problems, rewrite the solutions as either a terminating or nonterminating decimal.

OpenStudy (anonymous):

@jgirl128

OpenStudy (jgirl128):

I'm sorry, I'm not that good at math. try tagging some of the mathletes on here. sorry and good luck :)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8 @tkhunny

OpenStudy (tkhunny):

So? Do it! You should be able to do this all day. Billy had some eggs. When he went by Jane's house, she gave him half as many eggs as he already had. Little Stevie wasn't feeling well, so Billy gave him 1/3 of the eggs in his possession. When Billy got to the egg market, he had 30 eggs. How many eggs did Billy have when he started out? x + (1/2)x - (1/3)[x + (1/2)x] = 30 Fractions, Addition, and Subtraction. Done x + (1/2)x - (1/3)[x + (1/2)x] = 30 1.5x - (1/3)(1.5x) = 30 1.5x - 0.5x = 30 x = 30 Seriously, just make it up and make it consistent. For example, the first time I was just throwing in numbers, it appears that Mary gave him 8.5 eggs. That's no good. It has to be consistent all the way through!

OpenStudy (anonymous):

thanks my sister needs help with it.

OpenStudy (tkhunny):

Thanks for what? You can't use my word problem. Make you own!